OpenStudy (anonymous):

A parallelogram has a base that measures 5 units and a height that measures (x + 6) units. The area of this parallelogram is 20 square units. What must be the value of x? Choose one answer. a. x=4 b. x=-2 c. x=5 d. x=15

OpenStudy (anonymous):

(5)(x + 6) = 20 So, 5x + 30 = 20 5x = -10 x = -2

OpenStudy (anonymous):

Oh, BTW we got (5)(x + 6) = 20 from b x h = A

OpenStudy (anonymous):

You seem to have a lot of parallelogram questions!

OpenStudy (anonymous):

yes i do lol